OpenStudy (anonymous):

a house cast a 56ft shadow. a girl standing nearby cast a 7.2ft shadow. her height is 5.4ft. what is the height of the house.

OpenStudy (anonymous):

Draw a picture first

OpenStudy (anonymous):

7.2/5.4=56/x X= 56(5.4)/7.2=42feet The house tall is 42 feet

OpenStudy (anonymous):

no picture needed .... just proportions

OpenStudy (anonymous):

Do a proportion, tjhen cross multiply.

OpenStudy (anonymous):